OpenStudy (anonymous):

write the equation of a line with the given information: m = 2 and (3,14)

OpenStudy (bahrom7893):

y - 14 = 2(x-3)

OpenStudy (bahrom7893):

y = 2x - 6 + 14 y = 2x + 8

OpenStudy (bahrom7893):

my last problem for today.. nite folks, my days start early and end late..

OpenStudy (anonymous):

Wow. You did that quickly. Thank you!

OpenStudy (bahrom7893):

i should've been asleep 2 hrs ago, dunno how ima wake up for class 2morrow..

OpenStudy (bahrom7893):

U are very welcome!

OpenStudy (anonymous):

I hope you're able to wake up! Good night. Thanks again for the quick help.